# User Simulation for Evaluating Information Access Systems

### v1

@@ 3.2.3 Modeling Search Strategies
Azzopardi:2011:SIGIR applies economic theory to study the trade-off between (number of queries issued) and assessing (number of results examined) in sessions. It is found that a more advanced retrieval model (BM25) supports a greater variety of search strategies, in order to achieve the same gain, compared to simpler models (like TF-IDF or Boolean retrieval). Baskaya:2012:SIGIR study the effectiveness of various query modification and result scanning strategies, given an overall time constraint. Depending on the device used for searching, desktop PC vs. smart phone, the costs associated with query formulation differ—as it takes more time to enter queries using a smart phone than using a regular keyboard. The authors show that the device and overall search time allocation profoundly affect what kind of interactive behaviour can be successful. Nonetheless, the more time available, the less it matters how one searches. The authors also show that traditional rank-based metrics that focus on the quality of ranking alone (such as MAP and NDCG) are insufficient or even misleading when input costs are taken into account.
Smucker:2016:CHIIR consider the scenario of a searcher switching from one search engine to another, in an expectation of better results. By examining the gain curves for a set of TREC queries, the authors identify four classes of switching behaviour: never switch (ranking A has consistently faster gain rate than ranking B), switch immediately (the opposite of never switch), moot switch (rankings A and B have similar performance, switching neither helps nor hurts), and complex switch (one ranking has a high rate of gain initially, but it is not sustained). It should be noted that while it is possible to calculate optimal switching behaviour, modeling this user decision would require “some user computable rate of gain and its comparison to an expected rate” [Smucker:2016:CHIIR].
@@
3.3 User Simulation for Recommender Systems and Conversational Assistants
@@
Compared with the study of user behaviour in search engines, much less work has been done on analyzing user behaviour in recommender systems or conversational assistants, which is likely due to the fact that the search engines have had a much longer history of commercial applications and deployment in the real world than recommender systems and conversational assistants, which have appeared more recently. However, it is worth noting that because search engines, recommender systems, and conversational assistants are all specific forms of information access systems, the users of those systems may also show some similar common behaviour or preferences. As such, the findings and techniques developed from research of user simulation in the context of search engines may also be applicable to recommender systems and/or conversational assistants. For example, recommendation and search are so closely related (cf. Section 2.1) that the way users interact with a recommended (ranked) list of items may be essentially the same as how they would interact with a search result list.
In the rest of this section, we briefly discuss the user simulation work done in the context of recommender systems and conversational assistants.
3.3.1 Recommender Systems
@@
Most of the existing work on analyzing user behaviour in recommender systems has focused on predicting whether a user would click on a recommended item (i.e., click modeling) often by using a broad spectrum of features computed for a candidate user-item pair. While many models are not interpretable, some work has suggested that a user’s decision on whether to click on a recommended item is influenced by the same user’s past actions and may also be correlated with the actions of other similar users. Some work has attempted to model consumer behaviour in an e-commerce context [Fleder:2007:EC, Fleder:2009:MS], analyze news recommenders (with respect to long-tail novelty and unexpectedness) [Bountouridis:2019:FAT], and measure gender bias in music recommenders [Ferraro:2021:CHIIR].
@@
Unlike the case of search engines, where simulation is typically concerned with the behaviour of individual users, simulations for recommender systems are usually performed with the goal of understanding how certain algorithmic choices impact a population of users and study different forms of bias, e.g., preference bias [Zhou:2021:SSRN] and popularity bias [Yao:2021:arXiv].
@@
These are only a few simulation studies on recommender systems with a focus on evaluating system performance in terms of accuracy or diversity. Szlvik:2011:ICWSM study different user choice models in terms of which users of a movie recommender system would choose from the list of recommendations. These include always accepting the recommendations (Yes-men), rating only movies that are globally the highest rated (Trend-followers), and making random selections (Randomizers). Their results show that (1) if the recommender system offers a personalized list of movies, then diversity is expected to be higher but overall ratings will suffer; (2) if only popular movies are recommended, then ratings will be higher but diversity will drop; and (3) if random movies are recommended then ratings are likely to become low, but users will receive a wide range of recommendations. Jannach:2015:UMUAI assess the effect of different recommendation strategies on the diversity of recommendations and find that the concentration bias of some algorithms on popular items can lead to a possibly undesired “blockbuster” effect.
@@
The problem of recommendation has so far been framed as mostly a single round of interaction (i.e., the system recommends information to a user), instead of an interactive session as in the case of a search system. Thus, there is a limited set of possible user actions to take, which makes a user’s behaviour less complicated than in the case of a search engine in some sense. Even when it is considered how the user responds to a recommended item, there appears to be no standard interface or interaction workflow assumed about a user’s further interactions, making the problem of user behaviour not well-defined. As future recommender systems become increasingly interactive, it is reasonable to anticipate more research on user simulation for recommender systems.
3.3.2 Conversational Assistants
@@
Conversational assistants represent an emerging and actively researched topic [Zamani:2023:arXiv]. There is already a clear trend of growth of research work in this area including work on user simulation. Building on research on task-oriented dialogue systems, Zhang:2020:KDD develop a user simulator specifically for the evaluation of conversational recommender systems use it to experimentally compare three preexisting conversational recommender systems. Zhang:2022b:SIGIR study how humans behave and formulate their utterances when a conversational assistant fails to understand them in order to mimic this behaviour in a simulator. Sekulic:2022:WSDM and Owoicho:2023:SIGIR use user simulation to facilitate the evaluation of mixed-initiative conversational search systems. Specifically, their simulators can answer clarification questions that a system may ask of its user.
@@
3.4 Applications of User Simulation
3.4.1 User Behaviour Analysis
Since a user simulator must be mathematically well defined, it serves as a formally specified hypothesis about a user’s behaviour. The benefit of a formally specified hypothesis is that the hypothesis can then be tested rigorously to examine its validity. This can be done by (1) designing focused user studies to examine to what extent the generated user behaviour of a formal user simulator match those of a real user or (2) by fitting a formal user model to the observed real user data such as a search log. While the user behaviour that can be studied in this way is limited to what formally can be modeled, this kind of research of user behaviour, which may be called “computational user modeling,” should presumably be our long-term goal in fully and precisely understanding a user. In this sense, advancement of user simulation generally should lead to improved understanding of users.
3.4.2 Evaluation of Interactive Systems
Second, virtually all the existing test collections (for both IR and recommender systems) can be regarded as an evaluation method based on some kind of simplified user simulators (see Chapter 4 for more discussion of this). A major limitation of these existing collections is that they cannot be used to evaluate an interactive system. Given the research progress already made in user simulation, we have the opportunity to address this limitation by systematically converting existing test collections into new collections with a wide range of user simulators trained based on the existing user data, thus enabling automatic evaluation of interactive search and recommender systems.
Finally, the application of user simulation for evaluation is closely related to its application to user studies in that in both cases interpretable user models are required in order for the applications to be successful. The requirement of interpretability in the case of user studies is obvious, but it is also required for evaluation for otherwise the evaluation results would not be so meaningful. With intrepretability, user simulation for evaluation would not only be meaningful but also be powerful in generating simulated users with different characteristics to allow for evaluation with different kinds of users. The integration of research of user studies and the development of formal interpretable user models is therefore highly desirable and may reinforce each other in accelerating research in both areas. Indeed, we envision a healthy ecosystem where user studies would lead to more realistic user simulators, which further enable the development of more effective interactive systems (e.g., with new functions and new interface features), leading to new opportunities of studying user behaviour in a new system environment (with new functions or new interfaces), and new generations of user simulators being developed.
3.4.3 Synthetic Data Generation
In addition to analysis of user behaviour and evaluation of systems, user simulation also has applications in synthetic data generation to create/augment training data for supervised machine learning, which tends to require large quantities of labeled examples of training data. Because manually annotating a data set is labor-intensive and costly, using user simulation to generate synthetic training data is appealing since it does not involve manual effort. Compared with other approaches to synthetic data generation, user simulation enables the generation of more interpretable synthetic data when the parameters used in the simulators are interpretable. Due to the scarcity of interaction data available (presumably because of issues such as the privacy), it is especially beneficial to leverage user simulation to generate interaction data, which can then be used for training reinforcement learning algorithms for learning an effective interaction policy. Work on generation of synthetic data includes, but not limited to (1) modeling clicks and interactions between search results [Zhang:2019:CIKM], (2) training spoken dialogue systems [Schatzmann:2006:KER, Schatzmann:2007:NAACL, Young:2010:CSL], (3) generating synthetic conversations by leveraging large pre-trained language models [Mohapatra:2021:EMNLP, Kim:2021:ACL, Wan:2022:EMNLP], (4) removing bias from RL-based training of recommender systems [Huang:2020:RecSys], and (5) deep reinforcement learning for IR [Zhang:2021:SIGIR] and RecSys [Chen:2021:arXiv].
@@ 3.4.3 Synthetic Data Generation
In this chapter, we discuss how user simulation can be used to evaluate information access systems. We begin with the observation that traditional evaluation measures used in information access may be viewed as naive user simulators. Next, we highlight the limitations of these metrics, and the additional components that would need to be modeled in order to going beyond them. We introduce a general evaluation framework for evaluating interactive information access systems using user simulation that can accommodate general interactions between a user and a system. As most existing work in this area has been done in the context of search engine applications, our review mostly covers such work. However, since most recommender systems present users with a ranked list of specific items from a collection as suggestions, most of the discussion is directly applicable to evaluation of ranked results from a recommender system as well.
@@ 4.1 Traditional Evaluation Measures as Naive Simulators
We discuss how traditional IR evaluation metrics may be viewed as evaluation using naive user simulators. This would provide some intuition of the general idea of simulation-based evaluation as well as motivate the need for a more general framework for simulating more realistic/sophisticated users.
@@
Early work on simulation also touched evaluation, but with no effort on simulating a user (see, e.g.,  [Cooper:1973:ISR, Gordon:1990:JASIS]). However, the expected search length can be regarded as an early discussion of how a metric can be viewed as quantifying a simulated user’s effort in achieving the goal of satisfying the information need [Cooper:1968:AD]. It has been recently shown that the whole Cranfield evaluation methodology can be regarded as evaluating an IR system using simulation of some kind of users [Zhang:2017:ICTIR].
The Cranfield evaluation methodology has so far been the standard way to evaluate information access sytems as well as virtually all the empirically defined tasks such as those in machine learning and natural language processing. The basic idea is to create a test collection that can be reused again and again for comparison of different systems (algorithms), enabling detailed study of individual components in a complex system by turning off/on each component. In the case of IR, the test collection generally consists of a collection of documents, a set of queries, and relevance judgments that indicate which documents in the collection are relevant/non-relevant to each query. To evaluate a system, we would run the system on a test collection to generate retrieval results for each query, and various evaluation metrics (e.g., Precision, Recall, Mean Average Precision, nDCG) would be used to quantify the performance of a system (as measured by the test collection).
@@
We now discuss how such a traditional evaluation methodology can be viewed as evaluation using naive user simulators. In general, we could interpret a measure such as Precision or Recall as the perceived utility of a retrieval result from the perspective of a simulated user. We can imagine the simulated user would sequentially “browse” the ranked list of retrieval results returned from the system and view every relevant document while skipping every non-relevant document. Now, suppose the user has viewed the top-$n$ documents, where there are $k$ relevant documents. In such a case, the user has spent effort to examine $N$ documents and gained relevant information from $k$ of them. Thus the Precision at $n$ documents, which is $k/n$, can be easily seen as measuring the ratio of the relevant information gain to the effort. When we fix $n$, Precision at $n$ documents can be interpreted measuring the gain of relevant information for a fixed budgeted effort. Similarly, Recall can be interpreted as measuring the progress toward completion of the task of finding all the relevant documents. The work [Zhang:2017:ICTIR] provides a detailed discussion of how many standard metrics can be interpreted with a user simulation framework.
@@
A particularly interesting observation is that the Average Precision can be interpreted as simulating a set of users with variable recall expectations. Another interesting observation is that the discounting coefficients in nDCG can be interpreted as the stopping probability at different positions, and the value of nDCG can be interpreted as the expected gain with consideration of the uncertainty of stopping at different positions. Here, we can see clearly the benefit of using a user simulation framework to examine a traditional measure. Consider when we use a measure such as NDCG@k for evaluation, from user simulation perspective, the parameter $k$ reflects where a user is expected to stop on a ranked list of results. Generally, this should depend on the information need behind the query, since an exploratory query with high recall may tend to be associated with a larger $k$, whereas known item search should use a smaller $k$. This means that ideally we should be using a query-specific value for $k$ when we evaluate a system over a set of queries. The argument for using variable $k$ for NDCG was indeed made recently by Karmaker:2020:CIKM, but embracing user simulation for evaluation would naturally reveal such a need.
@@
The time-based nDCG measure [Smucker:2012:SIGIR] can also be interpreted as modeling the probability of stopping more accurately based on the time spent on examining documents. In the next section, we will provide a more complete discussion of many measures that go beyond the traditional basic measures such as Precision and Recall; they generally incorporate more explicit modeling of user interactions and are thus even closer to evaluation using user simulation.
@@ 4.2 Going Beyond Traditional Evaluation Measures
Next, we discuss what other metrics need to be considered (going beyond the current metrics), especially those related to interactions with users, thereby motivating the various components that need to be modeled in a simulation framework, including modeling a user’s behaviour, modeling a user’s cost/effort, and modeling the reward (gain of relevant information). We will also discuss the need for modeling the variation of users as well as variable behaviours of the same user in different contexts. This would naturally motivate the need for a more general simulation framework to be discussed in detail in the next section.
@@ 4.2.1 Measures based on Explicit Models of User Behaviour
The Rank-Biased Precision (RBP) measure [Moffat:2008:TOIS] makes an explicit assumption about the user’s behaviour, which is to continue examining the next document with a probability of $p$. This provides a way to estimate the probability that a document at a particular rank position $i$ would be examined by a user: $p^{i-1}$ with the assumption that the very first one would be always examined. Rank-Biased Precision is then defined as to measure the expected utility to a user with this kind of behaviour:
$RBP(r_{1},...,r_{d})=(1-p).\sum_{i=1}^{d}r_{i}.p^{i-1}~{},$ (4.1)
where $r_{i}$ is the gain (reward) of relevance information as defined in the nDCG measure and can be flexibly defined based on an application need. There is no explicit modeling of effort, but it can be inferred as uniform cost in examining any document. Unlike the regular Precision measure, where the user is assumed to have a fixed budget for examining documents, RBP assumes that a user has a fixed level of patience in examining the results. In this measure, the task is not explicitly defined, but it is related to the “patience parameter” $p$ in the sense that a user who wants higher recall can be assumed to have a higher $p$.
@@
A more general formal framework for model-based measures was proposed in [Carterette:2011:SIGIR], where a measure is formalized as quantifying the accumulated utility over all the documents examined by a user who adopts a certain browsing model, thus introducing explicitly three component models: (1) a browsing model that captures how a user interacts with the ranked list of results, (2) a document utility model that captures the utility of a single document, and (3) a utility accumulation model that captures how the utility of each individual document would be combined to capture the total utility of all the examined documents. The separation enables different ways of instantiating each component model and combination of those different instantiations to obtain many different specific measures. In [Carterette:2011:SIGIR], it was shown that the framework can not only cover many existing measures and provide an interpretation of those measures to reveal the assumed component models used in each of them, but also suggest many interesting new measures that might also be meaningful. Although the modeling of cost was not explicitly included in this framework, conceptually, this framework has paved way for further generalization of model-based evaluation measures to the general user simulation-based evaluation methodology advocated in [Zhang:2017:ICTIR], where the browsing model would be generalized to a more general user simulator that can simulate not only browsing behaviour over a ranked list, but also any user action, including, e.g., query formulation/reformulation. We will describe this general framework later in this chapter.
The early IR measures have generally defined reward based on relevance-based gains. Later, novelty and diversity of the search results were also considered as part of the reward in some extended measures that consider both relevance and novelty, e.g., Normalized Discounted Cumulative Utility (NDCU) [Yang:2007:SIGIR] and $\alpha$-NDCG [Clarke:2008:SIGIR]. From the perspective of user simulation, their main differences are in the improvement of the modeling of reward as perceived by the simulated user.

### v2

@@ 3.3.3 Modeling Search Strategies
Azzopardi:2011:SIGIR applies economic theory to study the trade-off between querying (number of queries issued) and assessing (number of results examined) in sessions. It is found that a more advanced retrieval model (BM25) supports a greater variety of search strategies, in order to achieve the same gain, compared to simpler models (like TF-IDF or Boolean retrieval). Baskaya:2012:SIGIR study the effectiveness of various query modification and result scanning strategies, given an overall time constraint. Depending on the device used for searching, desktop PC vs. smart phone, the costs associated with query formulation differ—as it takes more time to enter queries using a smart phone than using a regular keyboard. The authors show that the device and overall search time allocation profoundly affect what kind of interactive behaviour can be successful. Nonetheless, the more time available, the less it matters how one searches. The authors also show that traditional rank-based metrics that focus on the quality of ranking alone (such as MAP and NDCG) are insufficient or even misleading when input costs are taken into account.
Search engine switching has been studied empirically in [White:2009:CIKM, Guo:2011:SIGIR]. Smucker:2016:CHIIR consider the scenario of a searcher switching from one search engine to another, in an expectation of better results. By examining the gain curves for a set of TREC queries, the authors identify four classes of switching behaviour: never switch (ranking A has consistently faster gain rate than ranking B), switch immediately (the opposite of never switch), moot switch (rankings A and B have similar performance, switching neither helps nor hurts), and complex switch (one ranking has a high rate of gain initially, but it is not sustained). It should be noted that while it is possible to calculate optimal switching behaviour, modeling this user decision would require “some user computable rate of gain and its comparison to an expected rate” [Smucker:2016:CHIIR].
@@
3.4 User Simulation for Recommender Systems
@@
Compared with the study of user behaviour in search engines, much less work has been done on analyzing user behaviour in recommender systems. However, since search engines and recommender systems are closely related (cf. Section 2.1), the findings and techniques developed from research of user simulation in the context of search engines may also be applicable to recommender systems. For example, the way users interact with a recommended (ranked) list of items may be essentially the same as how they would interact with a search result list.
@@
Most of the existing work on analyzing user behaviour in recommender systems has focused on predicting whether a user would click on a recommended item (i.e., click modeling) often by using a broad spectrum of features computed for a candidate user-item pair. While many models are not interpretable, some work has suggested that a user’s decision on whether to click on a recommended item is influenced by the same user’s past actions and may also be correlated with the actions of other similar users. Some work has attempted to model consumer behaviour in an e-commerce context [Fleder:2007:EC, Fleder:2009:MS, Zhang:2020:ISR, Chaney:2021:arXiv], analyze news recommenders (with respect to long-tail novelty and unexpectedness) [Bountouridis:2019:FAT], and measure gender bias in music recommenders [Ferraro:2021:CHIIR].
@@
Unlike the case of search engines, where simulation is typically concerned with the behaviour of individual users, simulations for recommender systems are usually performed with the goal of understanding how certain algorithmic choices impact a population of users and study different forms of bias, e.g., preference bias [Zhou:2021:SSRN] and popularity bias [Yao:2021:arXiv], or long-term dynamics [Hazrati:2024:UMUAI, Yao:2023:ICML].
@@
These are only a few simulation studies on recommender systems with a focus on evaluating system performance in terms of accuracy or diversity. Szlvik:2011:ICWSM study different user choice models in terms of which users of a movie recommender system would choose from the list of recommendations. These include always accepting the recommendations (Yes-men), rating only movies that are globally the highest rated (Trend-followers), and making random selections (Randomizers). Their results show that (1) if the recommender system offers a personalized list of movies, then diversity is expected to be higher but overall ratings will suffer; (2) if only popular movies are recommended, then ratings will be higher but diversity will drop; and (3) if random movies are recommended then ratings are likely to become low, but users will receive a wide range of recommendations. Jannach:2015:UMUAI assess the effect of different recommendation strategies on the diversity of recommendations and find that the concentration bias of some algorithms on popular items can lead to a possibly undesired “blockbuster” effect. In recent work, Rahdari:2024:TORS demonstrate with the help of simulation that “users find their desired item more efficiently when recommendations are presented as a list of carousels compared to a simple ranked list.”
@@
The problem of recommendation has so far been framed as mostly a single round of interaction (i.e., the system recommends information to a user), instead of an interactive session as in the case of a search system. Thus, there is a limited set of possible user actions to take, which makes a user’s behaviour less complicated than in the case of a search engine in some sense. Simulations have in fact been used in interactive recommender systems for evaluating critiquing and comparison-based feedback approaches [McGinty:2011:bookchapter]. However, there appears to be no standard interface or interaction workflow assumed about a user’s further interactions, making it challenging to study user behaviour. As future recommender systems become increasingly interactive, it is reasonable to anticipate more research on user simulation for recommender systems.
3.5 User Simulation for Conversational Assistants
Conversational assistants represent an emerging and actively researched topic [Zamani:2023:FnTIR]. There is already a clear trend of growth of research work in this area including work on user simulation, especially in the context of task-oriented dialogue systems [ElAsri:2016:Interspeech, Gur:2018:SLT, Gur:2018:SLT, Lin:2021:SIGDIAL, Kreyssig:2018:SIGDIAL, Lin:2022:SIGDIAL, Hu:2023:CIKM, Davidson:2023:arXiv].
@@ 3.5 User Simulation for Conversational Assistants
Building on research on task-oriented dialogue systems, Zhang:2020:KDD develop a user simulator specifically for the evaluation of conversational recommender systems use it to experimentally compare three preexisting conversational recommender systems. Zhang:2022b:SIGIR study how humans behave and formulate their utterances when a conversational assistant fails to understand them in order to mimic this behaviour in a simulator.
Another line of work utilizes large language models given their ability to generate fluent, human-like responses to a variety of inputs in natural language. Sekulic:2022:WSDM and Owoicho:2023:SIGIR use user simulation to facilitate the evaluation of mixed-initiative conversational search systems. Specifically, their simulators can answer clarification questions that a system may ask of its user as well as provide explicit feedback on the presented results.
@@
In this chapter, we discuss how user simulation can be used to evaluate information access systems. We begin with the observation that traditional evaluation measures used in information access may be viewed as naive user simulators. Next, we highlight the limitations of these metrics, and the additional components that would need to be modeled in order to going beyond them. We then introduce a general evaluation framework for evaluating interactive information access systems using user simulation that can accommodate general interactions between a user and a system. While our discussion is mostly done in the context of search engine applications, the main ideas, frameworks, and methods are all generally applicable to recommender systems and conversational assistants as well.
@@ 4.1 Traditional Evaluation Measures as Naive Simulators
In this section, we discuss how traditional IR evaluation metrics may be viewed as evaluation using naive user simulators. This would provide some intuition of the general idea of simulation-based evaluation as well as motivate the need for a more general framework for simulating more realistic/sophisticated users.
@@
Early work on simulation in IR already touched evaluation (see, e.g.,  [Cooper:1973:ISR, Gordon:1990:JASIS]); for example, the expected search length can be regarded as an early discussion of how a metric can be viewed as quantifying a simulated user’s effort in achieving the goal of satisfying the information need [Cooper:1968:AD]. Interpretation of commonly used retrieval measures form user simulation perspective has been studied in  [Moffat:2008:TOIS], where a measure explicitly defined based on user modeling, called Rank-Biased Precision was proposed. Zhang:2017:ICTIR further show that the whole Cranfield evaluation methodology can be regarded as a special case of using user simulation for evaluation and all the major past IR measures can and should be interpreted in the framework of user simulation, where, e.g., Average Precision is shown to model the perceived utility by a group of users with variable recall requirements.
@@
In general, we could interpret a measure such as Precision or Recall as the perceived utility of a retrieval result from the perspective of a simulated user. We can imagine the simulated (naive) user would sequentially “browse” the ranked list of retrieval results returned from the system and view every relevant document while skipping every non-relevant document. Now, suppose the user has viewed the top-$n$ documents and found that $k$ of these are relevant. In such a case, the user has spent effort to examine $n$ documents and gained relevant information from $k$ of them. Thus the Precision at $n$ documents, which is $k/n$, can be easily seen as measuring the ratio of the relevant information gain to the effort. When we fix $n$, Precision at $n$ documents can be interpreted measuring the gain of relevant information for a fixed budgeted effort. Similarly, Recall can be interpreted as measuring the progress toward completion of the task of finding all the relevant documents. Zhang:2017:ICTIR provide a detailed discussion of how multiple standard metrics can all be interpreted with a user simulation framework.
@@
A particularly interesting observation is that Average Precision can be interpreted as simulating a set of users with variable recall expectations. Another interesting observation is that the discounting coefficients in NDCG can be interpreted as the stopping probability at different positions, and the value of NDCG can be interpreted as the expected gain with consideration of the uncertainty of stopping at different positions. Here, we can see clearly the benefit of using a user simulation framework to examine a traditional measure. Consider when we use a measure such as NDCG@k for evaluation, from an user simulation perspective, the parameter $k$ reflects where a user is expected to stop on a ranked list of results. Generally, this should depend on the amount of relevant information that the use rexpects to find, which often depends on the information need behind the query. For example, a query with high recall (e.g., from a user doing a literature survey) may tend to be associated with a larger $k$, whereas known item search should presumably use a smaller $k$. This means that ideally we should be using a query-specific value for $k$ when we evaluate a system over a set of queries. The argument for using variable $k$ for NDCG was indeed made recently by Karmaker:2020:CIKM, but embracing user simulation for evaluation would naturally reveal such a need.
@@
The time-based NDCG measure [Smucker:2012:SIGIR] can also be interpreted as modeling the probability of stopping more accurately based on the time spent on examining documents. In the next section, we will provide a more complete discussion of many measures that go beyond the traditional basic measures such as Precision and Recall; they generally incorporate more explicit modeling of user interactions and correspond to a more realistic model of user behaviour, and are thus even closer to evaluation using user simulation.
@@ 4.2 Going Beyond Traditional Evaluation Measures
We now discuss what additional metrics need to be considered (going beyond the current metrics), especially those related to interactions with users, thereby motivating the various components that need to be modeled in a simulation framework, including modeling a user’s behaviour, modeling a user’s cost/effort, and modeling the reward (gain of relevant information). We will also discuss the need for modeling the variation of users as well as variable behaviours of the same user in different contexts. This would naturally motivate the need for a more general simulation framework to be discussed in detail in the next section.
@@ 4.2.1 Measures Based on Explicit Models of User Behaviour
The Rank-Biased Precision (RBP) measure [Moffat:2008:TOIS] makes an explicit assumption about the user’s behaviour, which is to continue examining the next document with a probability of $p$. This provides a way to estimate the probability that a document at a particular rank position $i$ would be examined by a user is $p^{i-1}$, with the assumption that the very first one would be always examined. Rank-Biased Precision is then defined as to measure the expected utility to a user with this kind of behaviour:
$RBP(r_{1},...,r_{d})=(1-p)\sum_{i=1}^{d}r_{i}\,p^{i-1}~{},$ (4.1)
where $r_{i}$ is the gain (reward) of relevance information as defined in the NDCG measure and can be flexibly defined based on an application need. There is no explicit modeling of effort, but it can be inferred as uniform cost in examining any document. Unlike the classic Precision measure, where the user is assumed to have a fixed budget for examining documents, RBP assumes that a user has a fixed level of persistence in examining the results. In this measure, the task is not explicitly defined, but it is related to the “persistence parameter” $p$ in the sense that a user who wants higher recall can be assumed to have a higher $p$.
@@
A more general high-level formal framework for model-based measures was proposed in [Carterette:2011:SIGIR], where a measure is formalized as quantifying the accumulated utility over all the documents examined by a user who adopts a certain browsing model, thus introducing explicitly three component models: (1) a browsing model that captures how a user interacts with the ranked list of results, (2) a document utility model that captures the utility of a single document, and (3) a utility accumulation model that captures how the utility of each individual document would be combined to capture the total utility of all the examined documents. The separation enables different ways of instantiating each component model and combination of those different instantiations to obtain many different specific measures. In [Carterette:2011:SIGIR], it was shown that the framework can not only cover many existing measures and provide an interpretation of those measures to reveal the assumed component models used in each of them, but also suggest many interesting new measures that might also be meaningful. Although the modeling of cost was not explicitly included in this framework, conceptually, this framework has paved way for further generalization of model-based evaluation measures to the general user simulation-based evaluation methodology advocated in [Zhang:2017:ICTIR], where the browsing model would be generalized to a more general user simulator that can simulate not only browsing behaviour over a ranked list, but also any user action, including, e.g., query formulation/reformulation. We will describe this general framework later in this chapter.
The early IR measures have generally defined reward in terms of relevance-based gains. Later, novelty and diversity of the search results were also considered as part of the reward in some extended measures that consider both relevance and novelty, e.g., Normalized Discounted Cumulative Utility (NDCU) [Yang:2007:SIGIR] and $\alpha$-NDCG [Clarke:2008:SIGIR]. From the perspective of user simulation, their main differences are in the improvement of the modeling of reward as perceived by the simulated user. Specifically, NDCU measures gain not solely based on relevance, but also based on novelty as compared with the information already consumed by the user, and $\alpha$-NDCG evaluates search results based on topic-level gain, instead of document-level gain, thus redundant documents that cover similar topics would not be favoured.